OpenStudy (anonymous):

evaluate the expression for a=3 and b=3 a^3-2b^2=-1

OpenStudy (aric200):

3^3-2(3)^2=-1

OpenStudy (anonymous):

is that the answer

OpenStudy (aric200):

No that is me simply pluging in the numbers

OpenStudy (anonymous):

a^3-2b^2+1=0 plug in a=3,b=3 3^3-2(3)^2+1=0 27-18+1 =10